OpenStudy (iwanttogotostanford):

HELP PLEASE: Rewrite f(x) = 2(x − 1)2 + 3 from vertex form to standard form.

OpenStudy (anonymous):

the answer would be

OpenStudy (anonymous):

2x^2-5x+5=0

OpenStudy (iwanttogotostanford):

how may i ask? @huaman

OpenStudy (iwanttogotostanford):

these are the answer choices: f(x) = 2x2 + 5 f(x) = 4x2 − 8x + 7 f(x) = 2x2 + 4x + 5 f(x) = 2x2 − 4x + 5

OpenStudy (iwanttogotostanford):

@huaman

OpenStudy (anonymous):

yuse distribution. wait let me draw it

OpenStudy (iwanttogotostanford):

ok, remember the answer choices.:-)

OpenStudy (anonymous):

ohh sorry bro. i thought this was converting to stnadard form from linear functions. but maybe this website could help. sorry again bro. http://mathbitsnotebook.com/Algebra1/Quadratics/QDVertexForm.html

OpenStudy (iwanttogotostanford):

do you know the answer based on the choices? @huaman